OpenStudy (anonymous):

please help me find the values by referring to the graphs of the trigonometric functions. if the value is undefined write undefined. 1. tan(-3/2pi) 2. cot(3/2pi) 3. sec 4pi 4. csc(-7/2pi)

OpenStudy (solomonzelman):

Well knowing that pi =180 1. tan(-3/2pi) = tan(-270) 2. cot(3/2pi) = cot(270) 3. sec 4pi = sec(720) 4. csc(-7/2pi) = csc(-630) Is that it, or you need an exact value?

OpenStudy (anonymous):

i need the exact value

OpenStudy (solomonzelman):

you can use the trigonometric table, http://mathmistakes.info/facts/TrigFacts/learn/vals/sum.html

OpenStudy (solomonzelman):

you can use a circle to first reduce angles bigger than 360.

OpenStudy (solomonzelman):

tan(-3/2pi) = (as tan is negative in 2nd quadrant) and it is = -sqrt3

OpenStudy (anonymous):

sorry im a bit confused

OpenStudy (solomonzelman):

I wrote too many replies, lets work each one together.

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

Don't even know how to start, 1) It is going to be in second quadrant, because it's negative so,) lets find the positive value (we know from a trigonometric table, http://mathmistakes.info/facts/TrigFacts/learn/vals/sum.html that) tan(3/2 pi)= -sqrt3

OpenStudy (solomonzelman):

I am really bad at explaining this staff, I think you need a tutor from OS to do this, not someone like me....

OpenStudy (solomonzelman):

@agent0smith, I'm not good at explaining can you help?